OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x - 2) Answer -17 -27 17 27

OpenStudy (aroub):

27

OpenStudy (anonymous):

5(x+8)=7(x-2) 5x+40=7x-14 40+14=7x-5x 54=2x x=54/2 x=27